OpenStudy (danysanchy97):

Write 1,386 as the product of its prime factors. A. 2 x 3 x 3 x 7 x 11 B. 2 x 3 x 7 x 33 C. 2 x 693 D. 2 x 3 x 231

OpenStudy (danysanchy97):

@FortyTheRapper

OpenStudy (janu16):

1386 ÷ 2 = 693 693 ÷ 3 = 231 231 ÷ 3 = 77 77 ÷ 7 = 11 ∴ 1386 = 2 x 3 x 3 x 7 x 11 (answer A)

OpenStudy (janu16):

@danysanchy97

OpenStudy (littlet):

@Janu16 is correct.

OpenStudy (welshfella):

to find the prime factors keep dividing by prime numbers starting from 2 2 ) 1386 3 ) 693 3 ) 231 7 } 11 77

OpenStudy (welshfella):

the prime factors are on the left